1. Isaac pulls a 50 kg box containing the toy car as a gift of his father on his 10th birthday across a level floor with a force of 200 N. If the force acts at an angle of 30° up from the horizontal, and the coefficient of kinetic friction is 0.30, determine:

 

a. the normal force exerted on the box by the floor;

b. the frictional force exerted on the crate by the floor

c. the acceleration of the box.


1
Expert's answer
2021-05-14T18:22:58-0400

a)


"N=mg-F\sin{30}\\N=50(9.8)-200\sin{30}=390\ N"

b)


"F_f=\mu N=0.3(390)=120\ N"

c)


"ma=F\cos{30}-F_f\\50a=200\cos{30}-120\\a=1.1\frac{m}{s^2}"